OpenStudy (anonymous):

A polygon has 7 sides. What is the sum of the measure of the exterior angles of the polygon? 180° 900° 360° 1260°

OpenStudy (anonymous):

No matter what type of polygon you have, the sum of the exterior angles is ALWAYS equal to 360°.

OpenStudy (anonymous):

Two exterior angles can be formed at each vertex of a polygon. The exterior angle is formed by one side of the polygon and the extension of the adjacent side. For the hexagon shown at the left, <1 and <2 are exterior angles for that vertex.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Yup ^

OpenStudy (anonymous):

so is it 360?

OpenStudy (anonymous):

Sum exterior angles of any polygon = 360° (using one exterior angle at a vertex)

OpenStudy (anonymous):

It is 360.

OpenStudy (anonymous):

ohhhhhh!!! i see.. Thank you!!

OpenStudy (anonymous):

220 students were asked if they liked chocolate or vanilla ice cream. 140 said they liked chocolate and 120 said they liked vanilla. How many students liked both chocolate and vanilla? 260 100 80 40

OpenStudy (anonymous):

what is 220 + 140?

OpenStudy (anonymous):

@rosesmarie

OpenStudy (anonymous):

360

OpenStudy (anonymous):

wait i meant 120 +140

OpenStudy (anonymous):

250

OpenStudy (anonymous):

actually its 260

OpenStudy (anonymous):

now 260 - 220 = ?

OpenStudy (anonymous):

oh sorry being distracted

OpenStudy (anonymous):

40

OpenStudy (anonymous):

thats your answer

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

The division property of equality could be used to solve which of the following equations? 5x = 30 x + 3 = 7 (x + 2) (x - 2) = 0 = 16

OpenStudy (anonymous):

mind putting this in a new tab

OpenStudy (anonymous):

sure